OpenStudy (anonymous):

37 is what percent of 131? Round the answer to the nearest hundredth. My answer is 3.54% can someone check it for me? please.

OpenStudy (some.random.cool.kid):

hmm

OpenStudy (anonymous):

A percent can be found as: (x/y) * 100 Where x is the first number and y is the second. 37/131 = .2824 What percent is that when you multiply it by 100?

OpenStudy (some.random.cool.kid):

yeah i got 28.24

OpenStudy (anonymous):

Also, just as a rule when determining if a percent answer is valid: Check the quarters of the second number. 1/4 of 131 is 32.75. Does it seem right then if 37 accounts for only 3.5% of 131, even though it excedes the first quarter of 131?

OpenStudy (some.random.cool.kid):

"What number is 37 of 131?" we assume that you are asking what number is 37 percent of 131. Since percent means per hundred, you are basically asking how many is 37 per hundred of 131. To calculate 37 of 131, we solve the the problem by using the following formula: (X*Y)/100 where X=37 and Y=131. Thus, when we enter 37 and 131 into our formula, we get this: (37*131)/100 = answer First we multiplied 37 by 131 and then we divided that by 100. When we did this, we found the answer to be 48.47. Therefore, 37 percent of 131 is 48.47.

OpenStudy (some.random.cool.kid):

this is more accurate.

OpenStudy (some.random.cool.kid):

To get the solution, we are looking for, we need to point out what we know. 1. We assume, that the number 131 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 100% equals 131, so we can write it down as 100%=131. 4. We know, that x% equals 37 of the output value, so we can write it down as x%=37. 5. Now we have two simple equations: 1) 100%=131 2) x%=37 where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 100%/x%=131/37 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for 37 is what percent of 131 100%/x%=131/37 (100/x)*x=(131/37)*x - we multiply both sides of the equation by x 100=3.5405405405405*x - we divide both sides of the equation by (3.5405405405405) to get x 100/3.5405405405405=x 28.24427480916=x x=28.24427480916 now we have: 37 is 28.24427480916% of 131

OpenStudy (some.random.cool.kid):

step by step

OpenStudy (some.random.cool.kid):

understand it now?

OpenStudy (some.random.cool.kid):

final answer is 28.24